OpenStudy (anonymous):

How do I solve the integral: \[\int\limits_{0}^{1}\sqrt{1-x^{2}}x^2dx\] Thanks in advance!

OpenStudy (amistre64):

\[\int\limits_{0}^{1}\sqrt{1-x^{2}}x^2dx\]

OpenStudy (amistre64):

... now it formats ..

OpenStudy (anonymous):

Thanks

OpenStudy (amistre64):

what methods have you tried? or know of?

OpenStudy (anonymous):

looks like a trig sub to me

OpenStudy (anonymous):

good morning sensei!

OpenStudy (amistre64):

i was considering a u sub ...

OpenStudy (amistre64):

Howdy sate

OpenStudy (anonymous):

Well I suppose I could try partial integration, I have a formula for integrals of the form \[\int\limits_{}^{}\sqrt{a-x^2}\] I just think it is to complicated to solve. I hoped for a good substitution!

OpenStudy (anonymous):

try \[x=\sin(u)\] and then you will get powers of sine to integrate. use reduction formula to finish it

OpenStudy (anonymous):

trust me it is set up for trig sub. \[u=\sin(x)\] \[du = \cos(x)dx\] \[\int \cos(u) \sin^2(u)\cos(u)du \] \[\int \sin^2(u)\cos^2(u)du \]

OpenStudy (amistre64):

just taking a stab at it .... spose u = 1-x^2, then du = -2x dx such that dx = du/-2x \[\int \cfrac{\sqrt{u}x^2}{-2x}du\] \[\int \cfrac{\sqrt{u}x}{-2}du\] but x = sqrt(1-u) soo \[\int \cfrac{\sqrt{u}\sqrt{1-u}}{-2}du\] \[\int \cfrac{\sqrt{u}\sqrt{1-u}}{-2}du\] ... prolly not gonna play nice this way is it :)

OpenStudy (amistre64):

trig it is ;) good call sate

OpenStudy (anonymous):

then one of those damn reduction formulas i forget. i can look in a calc book because \[\int\sin^2(u)\cos^2(u)du\] is probably even an example.

OpenStudy (anonymous):

Hey thanks! I solved it by satellite's substitution x=sinu! The result then is integral sin^2(u) cos^2(u) du = 1/32 (4 u-sin(4 u))+constant

OpenStudy (anonymous):

Thanks for the effort!

OpenStudy (anonymous):

good work. glad you did it and not me. i always forget how to do powers of sine and cosine. you have to use some reduction formula

OpenStudy (anonymous):

ok ajahangir but of course now you have to go back, yes? \[u=\sin^{-1}(x)\] and so on

OpenStudy (anonymous):

Oh, I always forget that step.. Thanks, suppose the answer will not look as pretty anymore

OpenStudy (anonymous):

Actually, I do not have to go back if I transform the boundaries as well, right? If the boundaries are 1 and 0 for x, then they are (pi)/2 and 0 for x=sinu, right? This would give the result: pi/16 !

OpenStudy (anonymous):

that is a better idea than going back, and your answer is right